# Serializing Java Objects in Plain Code

### v1

@@ 6 Related Work
Closely related to our work is CoDeSe, proposed by Gligoric et al. [16]. CoDeSe, short for COde-based DEserialization and SErialization, is an approach to serialize objects as executable Java code focusing on performance. While sharing the same idea of plain-code serialization, our work is very different in the outcome. CoDeSe heavily uses the Java reflection API and the Unsafe class [37] to reconstruct an object and set its fields, which can be considered as low-level reconstruction. On the contrary, our approach never uses reflection and unsafe calls, and only uses application-specific setters and methods that are visible in the deserialization scope. Consequently, it can be argued that the code is more readable.

### v3

@@ 6 Related Work
Closely related to our work is CoDeSe, proposed by Gligoric et al. [16]. CoDeSe, short for COde-based DEserialization and SErialization, is an approach to serialize objects as executable Java code focusing on performance. While sharing the same idea of plain-code serialization, our work is very different in the outcome. The objective for serialization with CoDeSe is fast deserialization, and to achieve this, it heavily invokes the Java reflection API and the Unsafe class [39] to reconstruct an object and set its fields. We invite the reader to peruse Figure 3b of [16], which presents the serialized representation of a red-black tree object produced by CoDeSe. The generated code includes calls to unsafe.allocateInstance(...) which directs the JVM to create the object in memory, as well as calls to unsafe.putInt(...) and unsafe.putObject(...) to manipulate its fields. CoDeSe can also generate C code to reconstruct objects, which is then processed through the Java Native Interface (JNI). We argue that this use of reflection and the JNI can be considered as low-level reconstruction. On the contrary, our approach never uses reflection and unsafe calls, and only uses application-specific constructors and methods that are visible in the deserialization scope. The focus of plain-code serialization with ProDJ is idiomacity and readability, as illustrated by the serialized objects embedded within the test presented in 11.